We have a total amount of outcomes at:
6+4+7 = 17
Thus, let us compute the chance of NOT picking the Popular Science magazine needed (since each event is independent of one another):
P(E1) P(E2) P(E3) = 16/17 15/16 14/15 * 13/14 = 16P4/17P4 = 13/17
Therefore the probability of picking the December 1st edition is:
1 - P(En) = 1 -13/17 = 4/17
The probability of picking the December 1st edition is 4/17.
I hope my answer has come to your help. Thank you for posting your question here in Brainly.